OpenStudy (anonymous):

Algebra Help Please: What is the equation of the line in standard form that passes through the point (–1, 4) and is parallel to the line y = 2x – 3?

OpenStudy (anonymous):

Answer Choices: x – 2y = 6 x + 2y = 6 2x – y = –6 2x + y = –6

OpenStudy (anonymous):

I only have 1 set of numbers though ??

OpenStudy (anonymous):

Use point-slope formula: \[ y-y_0=m(x-x_0) \]We know \(m=2\), so plug in the point.

OpenStudy (anonymous):

I tried that but my answer doesn't match up with any of the answer choices ?? I got y=2x+15 What am i doing wrong??

OpenStudy (anonymous):

Wait... is it 2x-y=-6

OpenStudy (anonymous):

Here: \[\begin{align} y-4&=2(x+1)\\ y-4&=2x+2\\ y&=2x+6\\ y-2x&=6 \end{align} \]